# Uni-ISP: Unifying the Learning of ISPs from Multiple Cameras

### v1

@@ Abstract
Modern end-to-end image signal processors (ISPs) can learn complex mappings from RAW/XYZ data to sRGB (or inverse), opening new possibilities in image processing. However, as the diversity of camera models continues to expand, developing and maintaining individual ISPs is not sustainable in the long term, which inherently lacks versatility, hindering the adaptability to multiple camera models. In this paper, we propose a novel pipeline, Uni-ISP, which unifies the learning of ISPs from multiple cameras, offering an accurate and versatile processor to multiple camera models. The core of Uni-ISP is leveraging device-aware embeddings through learning inverse/forward ISPs and its special training scheme. By doing so, Uni-ISP not only improves the performance of inverse/forward ISPs but also unlocks a variety of new applications inaccessible to existing learned ISPs. Moreover, since there is no dataset synchronously captured by multiple cameras for training, we construct a real-world 4K dataset, FiveCam, comprising more than 2,400 pairs of sRGB-RAW images synchronously captured by five smartphones. We conducted extensive experiments demonstrating Uni-ISP’s accuracy in inverse/forward ISPs (with improvements of +1.5dB/2.4dB PSNR), its versatility in enabling new applications, and its adaptability to new camera models.
@@
Figure 1: We propose Uni-ISP, a model that unifies the learning of inverse and forward ISP behaviors of multiple cameras simultaneously. By leveraging the shared characteristics across various camera ISPs, our method can achieve higher performance in inverse and forward ISP (A) compared to previously learned ISP methods tailored for only one camera separately. Meanwhile, the device-aware property of the Uni-ISP enables new cross-camera ISP applications for a learned ISP model, including photographic appearance transfer (B and C), inter/extrapolation (D), and zero-shot image forensics (E and F).
1 Introduction
@@
Image Signal Processor (ISP) transforms raw image data captured by camera sensors into viewable formats such as sRGB, playing a pivotal role in determining the visual quality of photographs [16]. Through the meticulous design of their ISPs, various camera brands have cultivated distinctive photographic appearances that resonate with diverse user preferences [9, 33]. For example, Apple’s smartphone cameras are celebrated for their sharp and distinctive Apple feel, while Leica cameras are esteemed for their glow and deep color tones, contributing to the iconic Leica look.
@@
Recently, neural networks have been used to approximate the entire ISP or a specific module, aka learned ISP, bringing two main benefits. 1) Performance Enhancement. The powerful representation capabilities of neural networks enable learned ISPs to perform challenging tasks, such as hallucinating detailed content in highlights and shadows [42]. 2) New Functionalities. Learned ISPs introduce new functionalities, such as inverse ISP [38, 1], which converts sRGB images back to RAW/XYZ space, offering greater flexibility and potential for raw-domain enhancement and further manipulation, such as deblurring [22], denoising [32, 5, 3], HDR photography [23, 42], etc. These innovations expand the applications and potential of learned ISPs. However, current methods still design and train ISPs for individual camera models, which might limit the synergistic benefits across different ISPs (see Section 4.1). Moreover, versatile models have shown advantages in various fields in low-level vision [31, 21], high-level vision [19], and multi-modality [41, 39, 35]. As the number of camera models grows, individual learned ISPs may also lack the versatility and adaptability for widespread use, potentially making it unsustainable in the long term.
@@
In this paper, we aim to unify the learning of ISPs from different cameras, which offers two direct advantages. First, it enhances visual quality by leveraging the synergies among ISPs from multiple cameras. Unified learning enables ISPs to understand the underlying commonalities and differences in the diverse data, leading to overall improved visual performance. Second, unified learning provides novel applications, surpassing the capabilities of existing learned ISPs limited to inverse and forward ISPs. The unified learning enables new uses such as photographic appearance transfer, interpolation, and extrapolation across diverse camera models. Additionally, it supports zero-shot image forensics based on the self-consistency of ISP behaviors, including image-level source camera identification and pixel-level image splice detection.
@@
However, learning a device-aware ISP model for multiple cameras is far from trivial, presenting several challenges. First, we observe that simply mixing training data from multiple cameras in current models [1, 40, 38, 18] can not bring satisfactory performances. Therefore, we propose Uni-ISP, a novel unified ISP model for multiple cameras that contains several optimizable device-aware embeddings to learn the ISPs of different cameras. These device-aware embeddings enable the model to capture the specific characteristics tailored to individual devices, while the shared backbones capture underlying commonalities.
@@
Second, current ISP datasets do not contain synchronized sRGB-Raw image pairs captured by multiple cameras. Although this data is not necessary to learn the individual ISP, it is critical for learning a unified ISP that supports synergistic benefits across different camera models and enables the development of new applications. To address this, we develop a synchronized camera array with five smartphones and construct a novel dataset, FiveCam, which consists of 2,464 synchronized high-quality sRGB-Raw paired images with 4K spatial resolution. The captured dataset has a wide range of scenarios, from landscape to close-up, and contains different lighting conditions, including both indoor and outdoor settings during day and night.
@@
Third, given the inevitable misalignment in synchronized photo pairs taken by different cameras, a robust alignment and training scheme is required. To tackle this challenge, we first roughly align the images using optical-flow-based methods, which will introduce frequency bias in wrapped images. Then, we design a frequency bias correction (FBC) loss to mitigate texture blur. Additionally, we introduce the self-/cross-camera training schemes to facilitate applications across the same/different camera models.
@@
With all these three designs, our Uni-ISP can be applied to a wide range of image tasks (Figure 1), such as photographic appearance transfer, interpolation, and extrapolation across diverse camera models. Users can apply the visual characteristics of one camera model to another, achieving unique aesthetic effects. It also facilitates zero-shot image forensics tasks by utilizing the self-consistence of these ISP behaviors, including source camera identification and image splice detection. Extensive experiments show that Uni-ISP outperforms state-of-the-art methods with approximately 1.5dB PSNR in inverse ISP and approximately 2.4dB PSNR in forward ISP.
@@ 2 Related Work
Training neural networks to learn the inverse and forward processes of certain ISPs has been explored in recent research works [1, 40, 38]. CycleISP [40] maintains the self-contained property of inverse and forward ISP, where the cycle constraints are applied in addition to the independent supervision from the RGB or the raw ground truth. The InvISP [38] achieves the inverse and forward ISP in one self-contained invertible network. However, constraints that ensure the model’s invertibility also limit the network’s expression ability. ParamISP [18] utilizes the EXIF data in the model design. All existing methods ignore the variance of camera devices, which overlooks the potential value of the commonalities across different camera devices in the learned ISP task. In contrast, our Uni-ISP leverages these underlying commonalities by unifying learning of multiple cameras.
@@
Photorealistic image style transfer manipulates the aesthetic of an image, such as its color and tone, without distorting its original structure and content. Non-learning-based photorealistic image style transfer is primarily handled through traditional image processing techniques [29, 30, 27], including color grading [30], gamut optimization [27], etc.  Gharbi et al. [13] and  Xia et al. [37] propose learning-based bilateral grids for photorealistic image style transfer. The recent large generative models have also opened up new possibilities for photorealistic image style transfer [12, 20, 24, 4], but at a cost of huge computational resources and the risk of distorting the original content. For efficient and high-fidelity photorealistic image style transfer, NeuralPreset [17] is proposed for learning photorealistic image style transfer with a small memory footprint, which can be easily deployed on mobile devices. Unlike the aforementioned methods, our Uni-ISP learns the transfer of style-like visual feeling produced by certain camera ISPs, the photographic appearance, which is device-dependent and physically faithful.
@@
Image forensics focuses on validating and analyzing digital images to determine their origin by source camera identification or authenticity by splice detection. Source camera identification is highly related to the properties of the camera, which can be done by comparing the sensor pattern noise (SPN) and its main component photo response non-uniformity (PRNU) noise of a given image taken with an unknown camera and the reference image taken with known cameras.  Lukas et al. [25] extract the photo response non-uniformity (PRNU) using the discrete wavelet transform.  Chen et al. [7] propose a residual network for source camera identification.  Hui et al. [14] propose a multi-scale feature fusion network and a corresponding two-stage training scheme for this task based on the guidance of PRNU. Some methods are not based on the PRNU but on the lens distortions and auto white balance (AWB) algorithms [10, 2]. Unlike the existing methods that require specific training, our method achieves zero-shot source camera identification by implicitly identifying the whole ISP behavior. Our model can also perform image splice detection by slightly modifying the inference process for source camera identification.
3 Method
3.1 Overview
@@
Figure 2 shows the overview of Uni-ISP, which contains the inverse ISP module $g$ and forward ISP module $h$. Our model aims to be aware of various camera devices in the learned ISP tasks. Suppose the images we discussed here are all in size of $H\times W$ and channels of $C$. Given an sRGB image $I_{a}\in\mathbb{R}^{H\times W\times C}$ produced by the camera $a$, $\mathcal{E}_{a}\in\mathbb{R}^{D}$ represents the device-aware embedding for camera $a$ with size of $D$, the inverse ISP module $g$ learns to output the corresponding XYZ image $\hat{L}_{a}$ of camera $a$ given the input $I_{a}$:
@@
3.2 Model Design
As depicted in Figure 2, our Uni-ISP utilizes the inverse ISP module $g$ and forward ISP module $h$, each featuring an encoder-decoder architecture. Both modules incorporate Local Feature Extraction Blocks (LFEBs) for detailed local processing and Global Feature Manipulation Blocks (GFMBs) for broad image adjustments, mirroring the dual processing dynamics of real camera ISPs that manage both global operations like exposure compensation and color correction, and local tasks such as tone mapping and highlight recovery.
@@ 3.2 Model Design
Global ISP operations are significantly affected by parameters such as exposure time and ISO, which alter the overall appearance of the photos. Our GFMBs are designed to incorporate these parameters to modify the residual features from the encoder’s LFEBs. These manipulated features are then relayed to corresponding LFEBs in the decoder stage. Used camera parameters (exposure, ISO, and f-number) are extracted from the EXIF metadata of JPEG images produced by cameras.
@@
Figure 3: The illustration of frequency bias in dataset wrapped using optical flow method. The interpolation during the wrapping will make images look blurry compared to the original one, eliminating its high-frequency component.
3.3 Training Scheme
We craft a special training scheme for Uni-ISP, which consists of two types of training objectives, i.e., self-camera ISP objective and cross-camera ISP objective.
3.3.1 Self-Camera Training Objective
In the self-camera training objective, our Uni-ISP learns the inverse and forward ISP behavior of multiple cameras simultaneously.
@@ 3.3.1 Self-Camera Training Objective
3.3.2 Cross-Camera Training Objective
@@
where $\mathcal{E}_{b}$ indicate the device-aware embeddings of camera $b$. Here the $h$ takes the $\hat{L}_{a}$ input as input, which is different from the ground truth $L_{a}$ used in the self-camera training objective. The cross-camera training objective is defined to minimize the distance between $\hat{I}_{b}$ and the ground truth $I_{b}$.
In the above definition, the input image $I_{a}$ and the output ground truth $I_{b}$ can not be aligned since they are captured using different cameras. This makes pixel-level losses like L1-loss fail to drive the cross-camera training objective that minimizes the distance between $\hat{I}_{b}$ and $I_{b}$. Therefore, we first use the optical-flow-based method, RAFT [34], to wrap our dataset for the cross-camera training objective. All sRGB images $I_{b}$ taken with camera $b$ will be wrapped to be $I^{w}_{b}$, aligned with sRGB images $I_{a}$ taken with camera $a$. We apply this wrapping to every possible camera-to-camera sRGB pair in our dataset. Any areas that cannot be aligned will be annotated as occlusion and masked during training.
However, although the optical-flow-based wrapping method effectively aligns these images, it also introduces a frequency bias in our dataset. As demonstrated in Figure 3, the wrapped image tends to have less high-frequency details compared to the image before wrapping. If we use the aligned images to train our model directly, it will unexpectedly learn to smooth the images in the cross-camera ISP tasks.
To tackle this problem, we propose the frequency bias correction (FBC) loss for the cross-camera training objective. The Equation 9 shows the process of this task. The FBC loss can be written as:
@@
where $f_{low}$ is a low-pass filter and $\mathcal{L}_{Freq}$ is a loss in frequency domain. In our implementation, the $f_{low}$ is a Gaussian filter with a kernel size of 5, and we adopt the focal frequency loss [15] as the $\mathcal{L}_{Freq}$.
Finally, we use the FBC loss $\mathcal{L}_{FBC}$ to drive the cross-camera training. Detailed ablation results on this loss function are presented in Section 4.5.

### v2

# arXiv:2406.01003v2[cs.CV] 27 Feb 2026

@@ Abstract
Modern end-to-end image signal processors (ISPs) can learn complex mappings from RAW/XYZ data to sRGB (and vice versa), opening new possibilities in image processing. However, the growing diversity of camera models, particularly in mobile devices, renders the development of individual ISPs unsustainable due to their limited versatility and adaptability across varied camera systems. In this paper, we introduce Uni-ISP, a novel pipeline that unifies ISP learning for diverse mobile cameras, delivering a highly accurate and adaptable processor. The core of Uni-ISP is leveraging device-aware embeddings through learning forward/inverse ISPs and its special training scheme. By doing so, Uni-ISP not only improves the performance of forward and inverse ISPs but also unlocks new applications previously inaccessible to conventional learned ISPs. To support this work, we construct a real-world 4K dataset, FiveCam, comprising more than 2,400 pairs of sRGB-RAW images captured synchronously by five smartphone cameras. Extensive experiments validate Uni-ISP’s accuracy in learning forward and inverse ISPs (with improvements of +2.4dB/1.5dB PSNR), versatility in enabling new applications, and adaptability to new camera models.
@@
Figure 1: We propose Uni-ISP, a model that learns the inverse and forward ISP behaviors of multiple mobile cameras simultaneously. By leveraging the shared characteristics across various camera ISPs, our method can achieve higher performance in inverse and forward ISP (A) compared to previously learned ISP methods tailored for only one camera separately. Meanwhile, the device-aware property of the Uni-ISP enables new cross-camera ISP applications for a learned ISP model, including photographic appearance transfer (B and C), inter/extrapolation (D), and zero-shot image forensics (E and F).
I Introduction
@@
Image Signal Processor (ISP) transforms raw image data captured by camera sensors into viewable formats such as sRGB, playing a pivotal role in determining the visual quality of photographs [16]. Through carefully designed ISPs, camera brands have developed distinctive photographic styles that appeal to diverse user preferences [9, 33]. For example, Apple’s smartphone cameras are celebrated for their sharp and distinctive Apple feel, while Leica cameras are esteemed for their glow and deep color tones, contributing to the iconic Leica look.
@@
Recently, neural networks have been used to approximate the entire ISP or a specific module, aka learned ISP, bringing two main benefits. 1) Performance Enhancement. The powerful representation capabilities of neural networks enable learned ISPs to perform challenging tasks, such as hallucinating detailed content in highlights and shadows [43]. 2) New Functionalities. Learned ISPs introduce new functionalities, such as the inverse ISP [39, 1], which converts sRGB images back to RAW/XYZ space, offering greater flexibility and potential for raw-domain enhancement and further manipulation, such as deblurring [22], denoising [32, 5, 4], HDR photography [23, 43], etc. These innovations expand the applications and potential of learned ISPs. However, current methods design and train ISPs for individual camera models, which might limit the synergistic benefits across different ISPs (see Section IV-A). Moreover, versatile models have shown advantages in various fields in low-level vision [30, 21], high-level vision [19], and multi-modality [42, 40, 36]. As the number of camera models increases, individual learned ISPs may also lack the versatility and adaptability for widespread use, potentially making it unsustainable in the long term.
@@
In this paper, our aim is to build a model toward unifying the learning of ISPs from different mobile cameras, which offers two direct advantages. First, it improves visual quality by leveraging the synergies among ISPs from multiple mobile cameras. Unified learning enables ISPs to understand the underlying commonalities and differences in the diverse data, leading to improved visual performance. Second, unified learning provides novel applications, surpassing the capabilities of existing learned ISPs limited to inverse and forward ISPs. Unified learning enables new uses such as photographic appearance transfer, interpolation, and extrapolation across diverse camera models. Additionally, as a bonus, a unified ISP model has the potential to support zero-shot image forensics based on the self-consistency of ISP behaviors, including image-level source camera identification and pixel-level image splice detection.
@@
However, learning a device-aware ISP model for multiple cameras is far from trivial, presenting several challenges. First, we observe that simply mixing training data from multiple cameras in current models [1, 41, 39, 18] can not produce satisfactory performance. Therefore, we propose Uni-ISP, a novel unified ISP model for multiple cameras that contains several optimizable device-aware embeddings to learn the ISPs of different cameras. These device-aware embeddings enable the model to capture the specific characteristics tailored to individual devices, while the shared backbones capture underlying commonalities.
@@
Second, current ISP datasets do not contain synchronized sRGB-Raw image pairs captured by more than three cameras. Although these data are not necessary to learn the individual ISP, they are critical to learning a unified ISP that supports synergistic benefits across different camera models and enables the development of new applications. To address this, we develop a synchronized camera array with five smartphones and construct a novel dataset, FiveCam, which consists of 2,464 synchronized high-quality sRGB-Raw-paired images with 4K spatial resolution. The captured dataset has a wide range of scenarios, from landscape to close-up, and contains different lighting conditions, including both indoor and outdoor settings during day and night.
@@
Third, given the inevitable misalignment in synchronized photo pairs taken by different cameras, a robust alignment and training scheme is required. To tackle this challenge, we first roughly align the images using optical-flow-based methods, which will introduce frequency bias in warped images. Then, we design a frequency bias correction (FBC) loss to mitigate texture blur. Additionally, we introduce the self-/cross-camera training schemes to facilitate applications across the same/different camera models.
@@
With all of these three designs, our Uni-ISP can be applied to a wide range of image tasks (Figure 1), such as photographic appearance transfer, interpolation, and extrapolation across diverse cameras. Users can apply the visual characteristics of one camera model to another, achieving unique aesthetic effects. It also facilitates zero-shot image forensic tasks by utilizing the self-consistency of these ISP behaviors, including source camera identification and image splice detection. Extensive experiments show that Uni-ISP outperforms state-of-the-art methods with approximately 1.5dB PSNR in the inverse ISP and approximately 2.4dB PSNR in the forward ISP.
@@ II Related Work
Training neural networks to learn the inverse and forward processes of certain ISPs has been explored in recent research works [1, 41, 39]. CycleISP [41] maintains the self-contained property of inverse and forward ISP, where the cycle constraints are applied in addition to independent supervision from the RGB or the raw ground truth. InvISP [39] achieves the inverse and forward ISP in one self-contained invertible network. However, constraints that ensure the model’s invertibility also limit the network’s expression ability. ParamISP [18] uses the EXIF data in the model design. All existing methods ignore the variance of camera devices, overlooking the potential value of the commonalities across different camera devices in the learned ISP task. In contrast, our Uni-ISP leverages these underlying commonalities by unifying the learning of multiple cameras.
@@
Photorealistic image style transfer manipulates the aesthetics of an image, such as its color and tone, without distorting its original structure and content. Non-learning-based photorealistic image style transfer is mainly handled through traditional image processing techniques [28, 29, 27], including color grading [29], gamut optimization [27], etc.  Gharbi et al. [13] and  Xia et al. [38] propose learning-based bilateral grids for photorealistic image style transfer. Recent large generative models have also opened up new possibilities for photorealistic image style transfer [12, 20, 24, 3], but at the cost of huge computational resources and the risk of distorting the original content. For efficient and high-fidelity photorealistic image style transfer, NeuralPreset [17] is proposed for learning photorealistic image style transfer with a small memory footprint, which can be easily deployed on mobile devices. Unlike the aforementioned methods, our Uni-ISP learns the transfer of style-like visual feeling produced by certain camera ISPs, the photographic appearance, which is device-dependent and physically faithful.
@@
Image forensics focuses on validating and analyzing digital images to determine their origin by source camera identification or authenticity by splice detection. Source camera identification is highly related to the properties of the camera, which can be done by comparing the sensor pattern noise (SPN) and its main component photo response non-uniformity (PRNU) noise of a given image taken with an unknown camera and the reference image taken with known cameras.  Lukas et al. [25] extract the PRNU using the discrete wavelet transform.  Chen et al. [7] propose a residual network for source camera identification.  Hui et al. [14] propose a multi-scale feature fusion network and a corresponding two-stage training scheme for this task based on the guidance of PRNU. Some methods are not based on PRNU but on the lens distortions and auto white balance (AWB) algorithms [10, 2]. Unlike existing methods that require specific training, our method achieves zero-shot source camera identification by implicitly identifying the whole ISP behavior. Our model can also perform image splice detection by slightly modifying the inference process for source camera identification.
III Method
III-A Overview
@@
Figure 2 shows the overview of Uni-ISP, which contains the inverse ISP module $g$ and forward ISP module $h$. Our model aims to be aware of various camera devices in the learned ISP tasks. Assume all images discussed here have dimensions $H\times W$ and $C$ channels. Given an sRGB image $I_{a}\in\mathbb{R}^{H\times W\times C}$ produced by the camera $a$, $\mathcal{E}_{a}\in\mathbb{R}^{D}$ represents the device-aware embedding for camera $a$ with size of $D$, the inverse ISP module $g$ learns to output the corresponding XYZ image $\hat{L}_{a}$ of camera $a$ given the input $I_{a}$:
@@
III-B Model Design
As depicted in Figure 2, our Uni-ISP utilizes the inverse ISP module $g$ and forward ISP module $h$, each featuring an encoder-decoder architecture. Both modules incorporate Local Feature Extraction Blocks (LFEBs) for detailed local processing and Global Feature Manipulation Blocks (GFMBs) for broad image adjustments, mirroring the dual processing dynamics of real camera ISPs that manage both global operations like exposure compensation and color correction, and local tasks such as tone mapping and highlight recovery. The Device-aware Embedding Interaction Module (DEIM) adapts bottleneck features of the network into camera-specific features via an cross-attention mechanism with the learnable device-aware embeddings. Figure 3 shows the detailed structure of the three components of our model design.
Figure 3: Illustration of the sub-components (LFEB, GFMB, and DEIM) inside Uni-ISP, where IN refers to instance normalization, CA stands for the channel-wise image attention layer, SA stands for the spatial-wise image attention layer, and FC represents the fully-connected linear layer.
@@ III-B Model Design
ISP behaviors can be affected by different parameters of exposure time and ISO, especially our dataset is collected with built-in auto-exposure strategies of each cameras. Different ISP parameters can lead to different global appearance of the photos in the same scene, and these parameters can be easily accessed form the JPEG or Raw files output by cameras. Our GFMBs are designed to incorporate these parameters in the ISP learning process, which modifies the residual features from the encoder’s LFEBs. These manipulated features are then relayed to corresponding LFEBs in the decoder stage. Used camera parameters (exposure, ISO, and f-number) are extracted from the EXIF metadata of each JPEG images produced by cameras.
@@
Figure 4: The illustration of frequency bias in dataset warped using optical flow method. The interpolation during the warping will make images look blurry compared to the original one, eliminating its high-frequency component.
III-C Training Scheme
We craft a special training scheme for Uni-ISP, consisting of two types of training objectives: self-camera and cross-camera. Self-camera training ensures fair comparisons with prior single-camera ISP methods [18, 1, 39]. Cross-camera training enables more applications like photographic appearance transfer across devices.
III-C1 Self-Camera Training Objective
In the self-camera training objective, our Uni-ISP learns the inverse and forward ISP behavior of each camera itself simultaneously.
@@ III-C1 Self-Camera Training Objective
III-C2 Cross-Camera Training Objective
@@
where $\mathcal{E}_{b}$ indicate the device-aware embeddings of camera $b$. Here, $h$ uses $\hat{L}_{a}$ as input, which is different from the ground truth $L_{a}$ used in the self-camera training objective. The cross-camera training objective is defined to minimize the distance between $\hat{I}_{b}$ and the ground truth $I_{b}$.
In the above definition, the input image $I_{a}$ and the output ground truth $I_{b}$ are not aligned since they are captured using different cameras. This makes pixel-level losses such as L1-loss fail to drive the cross-camera training objective that minimizes the distance between $\hat{I}_{b}$ and $I_{b}$. Therefore, we first use the optical-flow-based method, RAFT [34], to warp our dataset for the cross-camera training objective. All sRGB images $I_{b}$ taken with camera $b$ will be warped to $I^{w}_{b}$, aligned with sRGB images $I_{a}$ taken with camera $a$. We apply this warping to every possible camera-to-camera sRGB pair in our dataset. Areas that cannot be aligned are marked as occlusions and masked during training.
However, although the optical-flow-based warping method effectively aligns these images, it also introduces a frequency bias in our dataset. As demonstrated in Figure 4, the warped image tends to have fewer high-frequency details compared to the image before warping. If we use the aligned images to train our model directly, it will unexpectedly learn to smooth the images in the cross-camera ISP tasks.
To address this problem, we propose the frequency bias correction (FBC) loss for the cross-camera training objective. The FBC loss mitigates the loss of high-frequency details due to optical-flow warping. It combines a low-frequency L1 part to ensure learning the low-frequent part with the warped image $I_{b}^{w}$ (e.g., the color and tones), and a frequency-domain part $\mathcal{L}_{\text{Freq}}$ to preserve the original image’s sharpness from $I_{b}$. The Equation 9 shows the process of this task. The FBC loss can be written as:
@@
where $f_{low}$ is a low-pass filter and $\mathcal{L}_{Freq}$ is a loss in the frequency domain. In our implementation, $f_{low}$ is a Gaussian filter with a kernel size of 5, and we adopt the focal frequency loss [15] as $\mathcal{L}_{Freq}$.
Finally, we use the FBC loss $\mathcal{L}_{FBC}$ to drive cross-camera training. Detailed ablation results on this loss function are presented in Section IV-E.